OpenStudy (hlpmeihlpu123):

so I'm in 9th grade and can barely multiply and cant divide i was wondering if you could give me some formulas for my notes so i can pass 9th grade

OpenStudy (divyankasc):

Like which lesson ?

OpenStudy (hlpmeihlpu123):

i forgot everything i learned last year and im already starting to fail

OpenStudy (divyankasc):

You need to work hard, a class 9 student can do 3 digit multiplications quickly

OpenStudy (divyankasc):

Because that's the basic. All formulas require that skills.

OpenStudy (hlpmeihlpu123):

well i was doing a dba and i didnt know how to find the area of a rug

OpenStudy (hlpmeihlpu123):

and i know most multiplication just some i dont

OpenStudy (hlpmeihlpu123):

i know how stupid i sound right now but its true

OpenStudy (divyankasc):

Uhm, can u plz explain what's DBA ?

OpenStudy (hlpmeihlpu123):

i take online school and its were i call my teacher and he ask me questions and i have to ask them i got caught using a caculator because i was taking to long

OpenStudy (hlpmeihlpu123):

i mean answer them not ask

OpenStudy (divyankasc):

So, well what help do u want now ?

OpenStudy (hlpmeihlpu123):

i just needed some basic formulas that will help me pass algebra

OpenStudy (hlpmeihlpu123):

the lessons are hard because im stupid and i dont know what im going to do this year

OpenStudy (divyankasc):

Like these ? (a + b)(a – b) = a2 – b2 (a + b + c)2 = a2 + b2 + c2 + 2(ab + bc + ca) (a ± b)2 = a2 + b2± 2ab (a + b + c + d)2 = a2 + b2 + c2 + d2 + 2(ab + ac + ad + bc + bd + cd) (a ± b)3 = a3 ± b3 ± 3ab(a ± b) (a ± b)(a2 + b2 m ab) = a3 ± b3 (a + b + c)(a2 + b2 + c2 -ab – bc – ca) = a3 + b3 + c3 – 3abc = 1/2 (a + b + c)[(a – b)2 + (b – c)2 + (c – a)2] when a + b + c = 0, a3 + b3 + c3 = 3abc

OpenStudy (hlpmeihlpu123):

yeah but like dumbed down to were i kow what they are which is going to be hard cause im pretty dumb